OpenStudy (kingkas):

Super confused: Two asteroids are 100,000 m apart. One has a mass of 3.5 106 kg. If the force of gravity between them is 1.05 10-4 N, what is the mass of the other asteroid? A. 4.1 108 kg B. 4.5 109 kg C. 4.1 109 kg D. 4.5 108 kg

OpenStudy (anonymous):

\[\huge F=G \frac{ m_1m_2 }{ r^2 }\] Where G=6.674×10−11 N⋅m^2/kg^2

OpenStudy (anonymous):

\[\huge m_2=\frac{ Fr^2 }{ Gm_1 }\]

OpenStudy (anonymous):

All the values you need on the right hand side are given to you in the problem